i have -3(4n+2)(n-6)=2(n+1)? find n to the nearest tenth thank you
Answer by CubeyThePenguin(3113)

-3(4n + 2)(n - 6) = 2(n + 1)
-3(4n^2 - 22n - 12) = 2n + 2
-12n^2 + 66n + 36 = 2n + 2
-12n^2 + 64n + 34 = 0

Use the quadratic formula: n = -0.5 and n = 5.8.